# SynthAorta: A 3D Mesh Dataset of Parametrized Physiological Healthy Aortas

### v1

@@ 3 Results
The dataset is stored efficiently since each mesh only needs to keep track of the $x,\,y\text{ and }z$ values of the nodes, which are stored in a standard csv file format. A single connectivity matrix is stored, as it is the same for every mesh. Moreover, we store connectivity matrices for $3$ coarser mesh levels in the same manner, with 224,1 792,2241792224,1\,792,224 , 1 792 , and 14 3361433614\,33614 336 elements, respectively. A simple function is provided as well, to extract the nodes for a coarser mesh into a new file.
@@ 4 Conclusion
The demo version of the dataset, with the code to convert the meshes to the well known .msh file format, is available on GitHub: https://github.com/domagoj-bosnjak/SynthAortaDemo. The full dataset will be made available post publication.
@@ Acknowledgements
The authors gratefully acknowledge Graz University of Technology, Austria for the financial support of the LEAD-project: Mechanics, Modelling and Simulation of Aortic Dissection.

### v2

@@ 3 Results
The dataset is stored efficiently since each mesh only needs to keep track of the $x,\,y\text{ and }z$ values of the nodes, which are stored in a standard csv file format. A single connectivity matrix is stored, as it is the same for every mesh. Moreover, we store connectivity matrices for $3$ coarser mesh levels in the same manner, with 224,1 792,2241792224,1\,792,224 , 1 792 , and 14 3361433614\,33614 336 elements, respectively. The main mesh contains 114 688114688114\,688114 688 elements. A simple function to extract the nodes for a coarser mesh into a new file is provided as well. For efficiency, the data is stored in binary format, using floats for 3D coordinates, and integers for connectivity matrices.
3.1 Mesh quality
@@
3.2 Aortic flow simulations
We first verify the suitability of the meshes through a convergence study. Three cardiac cycles $t\in(0,2.34~{}\mathrm{s}]$ are computed on successively finer grids of the grid with the worst element quality. Relative errors are computed taking the finest resolution as reference, see Tab. 3. We observe convergence of the flow rate and spatial mean pressure (exemplarily given for the outlet at abdominal level), and the WSS in the descending aorta. The WSS is subject to greater errors, since it is a function of the velocity gradient and the unit outward normal. Considering level $l=1$ for all meshes in what follows, we have thus strong indications that the quantities of interest (QoIs) are resolved with engineering precision (flow rates and pressure $<5\%$, WSS $<15\%$). However, similarly strict convergence studies for all cases in the SynthAorta dataset are computationally prohibitive. The resolution required depends on the physics of interest and can hence not be given in general, rendering the multiple grid resolutions provided in the SynthAorta dataset useful also in this sense.
Table 3: Relative errors in the flow rate $\epsilon_{Q}$ and spatial mean pressure $\epsilon_{\bar{p}}$ at the outlet at abdominal level, as well as the WSS $\epsilon_{\mathrm{WSS}}$ in the descending aorta under mesh refinement using mesh levels $l=0,1,2$.
Mesh level $\epsilon_{Q}$ [%] $\epsilon_{\bar{p}}$ [%] $\epsilon_{\mathrm{WSS}}$ [%]
0 20.54 3.61 99.81
1 1.91 0.22 11.00
2 (reference) - - -
Lastly, we provide hemodynamics QoIs for the entire SynthAorta dataset, where we reduce the interval of interest to the first cardiac cycle $t\in(0,0.78~{}\mathrm{s}]$, which is admissible given the tuned Windkessel parameters and omitting the first $t\approx 0.1~{}\mathrm{s}$, after which periodicity is already reached (discrepancies lie in the single-digit range comparing QoIs in the time interval [0.10, 0.78]0.100.78[0.10,\,0.78][ 0.10 , 0.78 ] with the periodic solution). In order to process massive amounts of fluid flow simulations of adequate resolution, optimized parameters in the simulation tools have to be fine-tuned. Inevitably, some simulations fail to complete if the settings are optimized for performance as necessary in the present case. Such, almost diverging, simulations yield unphysical results, and thus require physical filtering and outlier removal (for a detailed discussion of this dilemma see [59]). The corresponding results are depicted in Fig. 5 showing the QoI distributions in the SynthAorta dataset over the cardiac cycle. We observe that the deviation from the mean over the entire dataset is largest in the systolic phase, while the pressure and flow show significantly less variation than the WSS. This hints at the expected heavy impact of the shape hyperparameters on hemodynamics, while the relative importance of these parameters remain to be identified. Within this work, these results simply serve to demonstrate the applicability of the dataset to real-world physics. It lays the foundation for further investigations centered around sensitivity analysis and uncertainty quantification or model order reduction/surrogate modeling including machine learning techniques.
Figure 5: Temporal trace of the QoIs in the abdominal aorta: flow rate (left), mean pressure (middle) and spatial average of the wall shear stress (right) in the abdominal aorta. The deviation from the mean over the entire SynthAorta dataset (black dashed line) is largest in the systolic phase, while the pressure and flow show significantly less variation than the WSS.
@@ 4 Conclusion
Naturally, the approach comes with certain limitations. We do not claim that such a small parameter set, based on a single aortic model, captures the entire, comprehensive variety of the aortic morphology. More work needs to be invested to make it fully comprehensive. Extensions to certain pathological aortic geometries might be difficult due to the convolution surfaces’ inability to render sharp, intricate features. Independent parameter sampling can produce non-physiological geometries, an effect slightly exacerbated by the convolution surface blending, resulting in some examples being discarded. However, the detection of such situations is automated during the convolution surface construction. Finally, clinical validation of the resulting dataset is not in the scope of this paper.
@@
The full dataset is available open-source at the repository of the Graz University of Technology [71], containing the minimal code to load meshes and centerlines, as well as to export them to the well-known .msh format, enabling visualization and manipulation in the open-source tool Gmsh [50]. Since the dataset is fixed in the repository, any new developments or additions will be done on the corresponding Github project at https://github.com/domagoj-bosnjak/SynthAorta. Moreover, it contains $100$ examples from the dataset in the same format.
@@ Acknowledgements
The authors gratefully acknowledge the support by Graz University of Technology, Austria for the financial support of the LEAD-project: “Mechanics, Modelling and Simulation of Aortic Dissection”. R.S. was partially supported by the German research foundation through project no. 524455704 (“High-Performance Simulation Tools for Hemodynamics”) and partially supported by the EuroHPC joint undertaking Centre of Excellence dealii-X, grant agreement no. 101172493.